If x=1 and x= -8, then form a quadratic equation
Answer by mangopeeler07(462)

in a quadratic equation, for x, you solve for the value that would make an expression zero. So if x=1 and x= -8, just move backwards.

For positive x's
(x-b)=0
For negative x's
(x+b)=0

plug in your x's
(1-b)=0
(-8+b)=0

Solve for b's
b=1
b=8

So the original equation was
(x-1)(x+8)=0